# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v3

@@ Abstract
The forecasting of irregular multivariate time series (IMTS) is a critical task in domains like healthcare and climate science. However, this task faces two significant hurdles: 1) the inherent non-uniformity and missing data in IMTS complicate the modeling of temporal dynamics, and 2) existing methods often rely on computationally expensive architectures. To address these dual challenges, we introduce APN, a general and efficient forecasting framework. At the core of APN is a novel Time-Aware Patch Aggregation (TAPA) module that introduces an aggregation-based paradigm for adaptive patching, moving beyond the limitations of fixed-span segmentation and interpolation-based methods. TAPA first learns dynamic temporal boundaries to define data-driven segments. Crucially, instead of resampling or interpolating, it directly computes patch representations via a time-aware weighted aggregation of all raw observations, where weights are determined by each observation’s temporal relevance to the segment. This approach provides two key advantages: it preserves data fidelity by avoiding the introduction of artificial data points and ensures complete information coverage by design.The resulting regularized and information-rich patch representations enable the use of a lightweight query module for historical context aggregation and a simple MLP for final prediction. Extensive experiments on multiple real-world datasets demonstrate that APN establishes a new state-of-the-art, significantly outperforming existing methods in both prediction accuracy and computational efficiency.
@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao, Bi, and Chen 2018; Brouwer et al. 2019; Kidger et al. 2020; Ham, Kim, and Luo 2019; Vio, Díaz-Trigo, and Andreani 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and missing data (Yalavarthi et al. 2024; Zhang et al. 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al. 2024) and TimeCHEAT (Liu, Cao, and Chen 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
@@
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al. 2018; Rubanova, Chen, and Duvenaud 2019; Gravina et al. 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin 2021; Chen et al. 2018). GNNs-based models (Yalavarthi et al. 2024; Zhang et al. 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al. 2021); Transformer-based models (Chen et al. 2023; Zhang et al. 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al. 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al. 2023), SparseTSF (Lin et al. 2024b), and CycleNet (Lin et al. 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
To address the above challenges, we propose a general framework called APN. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b. Unlike methods that resample data points within a patch, TAPA directly aggregates the raw, irregular observations. It learns dynamically adjustable patch boundaries to flexibly frame regions of interest. Crucially, it then employs a time-aware soft-weighting mechanism to generate a patch representation, ensuring that all observations are holistically considered and contribute to the representation based on their temporal relevance to the patch. This approach not only adapts to local information density and captures complete semantic units but also inherently prevents information loss by design, as no data point is left uncovered. Based on these high-quality representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@
  • To address the IMTSF, we propose a general framework named APN. This framework leverages adaptive patching to generate high-quality and regular initial patch representations. Based on these representations, we employ a simple query module to integrate contextual information, enabling a lightweight yet powerful overall architecture.
  • We design a novel TAPA module to achieve adaptive patching. By learning dynamic patch boundaries and employing a time-aware soft-aggregation scheme over raw observations, TAPA produces high-quality regularized representations without interpolation. This mechanism effectively adapts to local variations in information density, captures complete semantics, and guarantees full coverage of all data points.
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
@@ 2 Related Work
2.1 Irregular Multivariate Time Series Forecasting
IMTSF is crucial for key domains such as healthcare and climate science. The inherent characteristics of IMTS, such as non-uniform sampling intervals and asynchronous channels, present significant challenges to regular time series forecasting models. To address these characteristics, researchers have proposed various IMTSF models. Some models employ approaches based on continuous-time dynamics (Chen et al. 2018; Rubanova, Chen, and Duvenaud 2019; Schirmer et al. 2022; Brouwer et al. 2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points. For instance, Neural Flows (Bilos et al. 2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers. GRU-ODE-Bayes (Brouwer et al. 2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism. Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al. 2024; Zhang et al. 2024; Liu, Cao, and Chen 2025) to capture complex dependencies in IMTS. For example, GraFITi (Yalavarthi et al. 2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs. T-PatchGNN (Zhang et al. 2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
The patch-based strategy, which has proven successful for regular time series (Nie et al. 2023; Wang et al. 2022), has been adapted for IMTS Forecasting (IMTSF). Initial adaptations, such as T-PatchGNN (Zhang et al. 2024) and TimeCHEAT (Liu, Cao, and Chen 2025), employed a fixed-span patching strategy. However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS. It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
To overcome this, adaptive patching emerged as a natural evolution. One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al. 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling. While effective for dense, regular data, this methodology is fundamentally flawed for IMTS. Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
In contrast, our work proposes a distinct aggregation-based adaptive patching paradigm, specifically designed for the challenges of IMTS. Instead of creating new, artificial data points, our TAPA module learns dynamic "soft windows" and computes each patch representation by performing a direct, weighted aggregation of all original observations. This strategy provides two critical advantages for IMTS:(1)Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation.(2)Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
@@ 3.1 Problem Definition
An Irregular Multivariate Time Series (IMTS) $\mathcal{O}$ consists of $N$ independent univariate sequences $\{o^{n}_{1:L_{n}}\}_{n=1}^{N}$. Each sequence $o^{n}_{1:L_{n}}$ comprises $L_{n}$ observations $(t_{i}^{n},v_{i}^{n})$, where the intervals between timestamps $t_{i}^{n}$ are irregular, and sampling across different variables is typically asynchronous. The IMTS forecasting task is defined as: given historical observations $\mathcal{O}$ and a set of prediction queries for all variables $\mathcal{Q}=\{[q_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$, construct and optimize a prediction model $\mathcal{F}(\cdot)$ capable of accurately predicting the future observed values $\hat{\mathcal{V}}=\{[\hat{V}_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$ corresponding to each query $q_{j}^{n}$, i.e.:
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
At the heart of APN lies the Time-Aware Patch Aggregation (TAPA) module, a novel mechanism designed to fundamentally reframe how we process irregular time series. Traditional methods rely on a paradigm of "hard segmentation," imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS. In stark contrast, TAPA introduces a "soft aggregation" paradigm. Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally-aware manner.
This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$. First, in Adaptive Patching, the model learns the dynamic temporal characteristics—the center and scale—of each patch’s field of influence. Second, in Weighted Aggregation, it computes a representation for each patch by performing a direct, weighted aggregation of all raw observations, where the weights are determined by each observation’s temporal relevance to the patch’s learned characteristics. This entire process transforms the irregular input sequence into a regularized, information-rich sequence of patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
@@
The first step is to define the temporal scope of each patch. Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$. This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data. For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed. Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
@@
$t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n},$ (2)
@@
where $T_{obs}$ is the time span of the historical observation window, and $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ represent the initial reference center and width, respectively. The use of $\exp(\lambda_{p}^{n})$ is a deliberate design choice, ensuring the learned window width remains strictly positive throughout optimization. This dynamic windowing mechanism allows the model to adaptively frame regions of interest, laying the foundation for a more meaningful and robust aggregation in the subsequent step.
@@
Enriching Observations with Temporal Context. To fully leverage temporal information, we first enrich each raw observation. A learnable time embedding, $TE(t_{i})\in\mathbb{R}^{D_{te}}$, is generated for each timestamp $t_{i}$. This embedding, composed of linear layers for scale and sine-activated layers for periodicity, captures both the absolute position and periodic patterns in time. It is then concatenated with the original value $v_{i}$ to form an augmented representation: $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
@@
$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right).$ (4)
Here, the first Sigmoid term models a smooth "falloff" from the right boundary, while the second models a smooth "rise" from the left boundary. Their product creates a continuous, bell-shaped weighting curve centered within the patch’s learned range. The term $\text{Softplus}(\kappa^{n})$ is a learnable, strictly positive temperature that controls the softness of the window edges. A smaller temperature leads to sharper, more defined boundaries, while a larger temperature creates gentler, more overlapping fields of influence, granting the model further adaptive flexibility.
A Crucial Design Choice for Information Integrity. The choice of this soft window function is critical. Since the Sigmoid function $\sigma(x)$ is strictly positive for all real inputs, the resulting weight $\alpha_{i,p}^{n}$ is also guaranteed to be positive for any observation $i$ and any patch $p$. This property is not a mere side effect; it is a fundamental design principle that mathematically guarantees complete information coverage. Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods where observations can be inadvertently discarded.
@@
3.4 Query-based Weighted Aggregation

### v4

@@ Abstract
The forecasting of irregular multivariate time series (IMTS) is crucial in key areas such as healthcare, biomechanics, climate science, and astronomy. However, achieving accurate and practical predictions is challenging due to two main factors. First, the inherent irregularity and data missingness in irregular time series make modeling difficult. Second, most existing methods are typically complex and resource-intensive. In this study, we propose a general framework called APN to address these challenges. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Additionally, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Experimental results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao2018; GRU-ODE; Neural-CDE; qiu2025DBLoss; qiu2025duet; wu2025k2vae; gao2025ssdts; hu2024multirc). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and missing data (GraFITi; tPatchGNN) in IMTS poses significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as tPatchGNN (tPatchGNN) and TimeCHEAT (TimeCHEAT), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is how to design an adaptive patching approach that can adapt to the local information density variations of irregular multivariate time series and capture complete semantic information.
@@
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural-ODE-based models (Neural-ODE; Latent-ODE; TG-ODE) require computationally intensive numerical solvers to accurately model continuous-time dynamics (mTAN; Neural-ODE). GNN-based models (GraFITi; tPatchGNN) are hindered by the overhead of complex graph construction and multi-round node information aggregation (BadGNN); Transformer-based models (ContiFormer; Warpformer), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (BadTransformer). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting tasks, models such as SparseTSF (lin2024sparsetsf), and CycleNet (lincyclenet) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is how to design an efficient model for IMTSF.
To address the above challenges, we propose a general framework called APN. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Based on the representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Results on multiple real-world datasets show that APN outperforms existing SOTA methods in both efficiency and accuracy.
@@
  • To address IMTSF, we propose a general framework named APN. This framework leverages adaptive patching to generate high-quality and regular initial patch representations. Based on these representations, we employ a simple query module to integrate contextual information, ensuring the effective design of the framework.
  • We design a novel TAPA module to achieve adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner, effectively adapting to local variations in information density and capturing complete semantics.
  • We conduct experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency. Additionally, all datasets and code are available at: https://github.com/decisionintelligence/APN.
@@ 2 Related Work
2.1 Progress in Irregular Multivariate Time Series Forecasting
IMTSF is crucial for key domains such as healthcare and climate science. The inherent characteristics of IMTS, such as non-uniform sampling intervals and asynchronous channels, present significant challenges to regular time series forecasting models. To address these characteristics, researchers have proposed various IMTSF models. Some models employ approaches based on continuous-time dynamics (Neural-ODE; Latent-ODE; CRU; GRU-ODE), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points. For instance, Neural Flows (Neural-Flows) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers. GRU-ODE-Bayes (GRU-ODE) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism. Other models leverage graph neural networks and attention mechanisms (GraFITi; tPatchGNN; TimeCHEAT) to capture complex dependencies in IMTS. For example, GraFITi (GraFITi) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs. tPatchGNN (tPatchGNN) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
The patch-based strategy, which has proven successful for regular time series (PatchTST; Crossformer; wu2025enhancing; qiu2025comprehensive; qiu2025dag; lu2025dtaf; wang2025plug), has been adapted for IMTS Forecasting (IMTSF). Initial adaptations, such as tPatchGNN (tPatchGNN) and TimeCHEAT (TimeCHEAT), employed a fixed-span patching strategy. However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS. It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
To overcome this, adaptive patching emerged as a natural evolution. One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (HDMixer)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling. While effective for dense, regular data, this methodology is fundamentally flawed for IMTS. Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
In contrast, our work proposes a distinct aggregation-based adaptive patching paradigm, specifically designed for the challenges of IMTS. Instead of creating new, artificial data points, our TAPA module learns dynamic “soft windows” and computes each patch representation by performing a direct, weighted aggregation of all original observations. This strategy provides two critical advantages for IMTS: (1) Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation. (2) Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
@@ 3.1 Problem Definition
An Irregular Multivariate Time Series (IMTS) $\mathcal{O}$ consists of $N$ univariate sequences $\{o^{n}_{1:L_{n}}\}_{n=1}^{N}$. Each sequence $o^{n}_{1:L_{n}}$ comprises $L_{n}$ observations $(t_{i}^{n},v_{i}^{n})$, where the intervals between timestamps $t_{i}^{n}$ are irregular, and sampling across different variables is typically asynchronous. The IMTS forecasting task is defined as: given historical observations $\mathcal{O}$ and a set of prediction queries for all variables $\mathcal{Q}=\{[q_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$, construct and optimize a prediction model $\mathcal{F}(\cdot)$ capable of accurately predicting future observed values $\hat{\mathcal{V}}=\{[\hat{v}_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$ corresponding to each query $q_{j}^{n}$, i.e.,
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
At the heart of APN lies the Time-Aware Patch Aggregation (TAPA) module, a novel mechanism designed to fundamentally reframe how we process irregular time series. Traditional methods rely on a paradigm of “hard segmentation,” imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS. In stark contrast, TAPA introduces a “soft aggregation” paradigm. Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally aware manner.
This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $o^{n}_{1:L_{n}}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$. First, in Adaptive Patching, the model learns the dynamic temporal characteristics—the center and scale—of each patch’s field of influence. Second, in Weighted Aggregation, it computes a representation for each patch by performing a direct, weighted aggregation of all raw observations, where the weights are determined by each observation’s temporal relevance to the patch’s learned characteristics. This entire process transforms the irregular input sequence into a regularized, information-rich sequence of patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
@@
The first step is to define the temporal scope of each patch. Instead of imposing predefined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$. This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data. For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed. Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
@@
$t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n}$ (2)
@@
where $T_{obs}$ is the time span of the historical observation window, with $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ representing the initial reference center and width, respectively. The use of $\exp(\lambda_{p}^{n})$ is a deliberate design choice, ensuring the learned window width remains strictly positive throughout optimization. This dynamic windowing mechanism allows the model to adaptively frame regions of interest, laying the foundation for a more meaningful and robust aggregation in the subsequent step.
@@
Enriching Observations with Temporal Context. To fully leverage temporal information, we first enrich each raw observation. A learnable time embedding $TE(t_{i}^{n})\in\mathbb{R}^{D_{te}}$ is generated for each timestamp $t_{i}^{n}$ in channel $n$. This embedding, composed of linear layers for scale and sine-activated layers for periodicity, captures both the absolute position and periodic patterns in time. It is then concatenated with the original value $v_{i}^{n}$ to form an augmented representation:
$\tilde{v}_{i}^{n}=\text{Concat}(v_{i}^{n},TE(t_{i}^{n}))$ (4)
@@
$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}^{n}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}^{n}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right)$ (5)
Here, the first Sigmoid term models a smooth “falloff” from the right boundary, while the second models a smooth “rise” from the left boundary. Their product creates a continuous, bell-shaped weighting curve centered within the patch’s learned range. The term $\text{Softplus}(\kappa^{n})$ is a learnable, strictly positive temperature that controls the softness of the window edges. A smaller temperature leads to sharper, more defined boundaries, while a larger temperature creates gentler, more overlapping fields of influence, granting the model further adaptive flexibility.
A Crucial Design Choice for Information Integrity. The choice of this soft window function is critical. Since the Sigmoid function $\sigma(x)$ is strictly positive for all real inputs, the resulting weight $\alpha_{i,p}^{n}$ is also guaranteed to be positive for any observation $i$ and any patch $p$. This property is not a mere side effect; it is a fundamental design principle that mathematically guarantees complete information coverage. Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods, where observations can be inadvertently discarded.
@@
3.4 Query-based Aggregation